### v1

@@ Abstract
Foundation models, pre-trained on large image datasets and capable of capturing rich feature representations, have recently shown potential for zero-shot image registration. However, their performance has mostly been tested in the context of rigid or less complex structures, such as the brain or abdominal organs, and it remains unclear whether these models can handle more challenging, deformable anatomy. Breast MRI registration is particularly difficult due to significant anatomical variation between patients, deformation caused by patient positioning, and the presence of thin and complex internal structure of fibroglandular tissue, where accurate alignment is crucial. Whether foundation model-based registration algorithms can address this level of complexity remains an open question. In this study, we provide a comprehensive evaluation of foundation model-based registration algorithms for breast MRI. We assess five pre-trained encoders, including DINO-v2, SAM, MedSAM, SSLSAM, and MedCLIP, across four key breast registration tasks that capture variations in different years and dates, sequences, modalities, and patient disease status (lesion versus no lesion). Our results show that foundation model-based algorithms such as SAM outperform traditional registration baselines for overall breast alignment, especially under large domain shifts, but struggle with capturing fine details of fibroglandular tissue. Interestingly, additional pre-training or fine-tuning on medical or breast-specific images in MedSAM and SSLSAM, does not improve registration performance and may even decrease it in some cases. Further work is needed to understand how domain-specific training influences registration and to explore targeted strategies that improve both global alignment and fine structure accuracy. We also publicly release our code at Github.
@@
Foundation model Breast MRI registration Zero-shot.
@@ 1 Introduction
Figure 1: Pipeline for Foundation Model-Based Breast Image Registration. The Found Encoder can be replaced with any pretrained vision foundation encoder.
In contrast, vision foundation models (VFMs) have gained attention for their strong generalization ability and zero-shot performance on downstream tasks such as segmentation and classification [24, 18, 9]. One of the few recent efforts, DINO-Reg [26], suggests that the rich semantic embeddings learned by DINO-v2 can be effectively applied to zero-shot registration in cross-domain medical images. Yet, it remains unclear how different pre-training strategies, such as masked autoencoders (MAE), vision-language models (VLMs), or medical domain-specific models like MedSAM, affect registration performance. Moreover, most studies focus on less deformable regions like abdominal organs with larger, more stable structures, while breast registration poses unique challenges with significant shape variation and thin, highly deformable tissue. How foundation model-based algorithms handle such complex anatomy remains poorly understood.
@@
    We conduct a comprehensive comparison of registration strategies according to the pretraining strategies that were applied to develop the foundation model, including contrastive learning, masked image modeling, and vision-language approaches, across both natural and medical imaging domains.
@@
    We design a series of breast registration tasks that reflect real-world challenges, covering longitudinal, cross-sequence, lesion-to-non-lesion, and PET/CT-to-MRI alignment.
@@
    We systematically evaluate the strengths and limitations of foundation model-based algorithms compared to other zero-shot approaches.
@@ 2 Method
2.1 Foundation model-based registration
Our training-free, foundation model-based registration pipeline, inspired by the framework of  [26], includes three main stages: (1) feature extraction, (2) dimension reduction, and (3) feature-based registration, as shown in Fig. 1.
@@ 2.1 Foundation model-based registration
To prepare for volumetric registration, we reshape $\tilde{F}_{\text{ref}}^{z}$, $\tilde{F}_{\text{mov}}^{z}$ back to a 3D spatial format. Specifically, each slice-level feature map $\tilde{F}^{z}\in\mathbb{R}^{N\times d}$ is first reshaped into a 2D grid of size $w\times h\times d$, where $w$ and $h$ represent the number of patch tokens along the width and height, respectively. These values depend on the encoder architecture, and $w\times h$ does not necessarily equal $N$, as some encoders include a global CLS token in $F^{z}$ that is excluded prior to reshaping. Stacking the reshaped slices along the $z$-axis results in a 4D volume of size $w\times h\times Z\times d$. This intermediate volume is then upsampled to match the original spatial resolution, resulting in a final feature volume of size $W\times H\times Z\times d$, which aligns with the input dimensions and serves as input to the registration step.
@@
Here, $\phi$ is the dense displacement field, $\text{Sim}(\cdot,\cdot)$ denotes a similarity metric (e.g., sum of squared differences or local cross-correlation), and $\lambda$ controls the smoothness constraint. The convex optimization is solved within a discrete search space to efficiently identify a globally optimal deformation.
@@ 2.2 Foundation models
SAM [13]: Trained on large-scale natural image segmentation datasets using a promptable segmentation objective. The model adopts a ViT-based encoder and was pre-trained with MAE, followed by additional prompt-based segmentation training. SAM has also been widely used for general vision tasks.
MedSAM [17]: A domain-specific version of SAM, fine-tuned on various medical image datasets including MRIs under a prompt-based setting to improve performance on medical image segmentation tasks.
SSLSAM [10]: Fine-tuned SAM on various MRI datasets without labels under the MAE setting. Noticeably, their pre-trained dataset also contains one publicly available breast dataset for dynamic pre-contrast MRIs [16].
@@
Feature extraction follows the input size and normalization required by each foundation model (Appendix 0.A). We also evaluate four non-foundation zero-shot models: Affine, SyN, NiftyReg, and ConvexAdam, detailed in Appendix 0.B.
@@ 3.1 Datasets preparation
Exam and mask collection: For the registration tasks, we collected both internal and external datasets including internal breast MRI exams and external PET-CT/MRI pairs from the BREAST-DIAGNOSIS collection [4]. We include patients with multiple time-stamped exams and pre-contrast axial views, and exclude volumes from patients who underwent mastectomy. For masks, we include three targets for our registration evaluation: breast, breast FGT, and organ. Breast and FGT masks are generated using a 3D V-Net-based model followed by manual correction, while organ and lesion masks are obtained using a nnU-Net-based segmentation model. Full preprocessing and annotation details for exams and masks are provided in Appendix 0.C and Appendix 0.D respectively.
@@
Tasks: In this study, we design four tasks to comprehensively evaluate the registration performance of multiple models. Task 1 involves registering breast MRI scans from different dates and years (longitiduanl exams) with the same image sequence (all images are dynamic pre-contrast). Task 2 focuses on registering longitiduanl breast MRI exams with different image sequences, where the fixed images are set to the dynamic pre-contrast sequence. The category of moving image sequences is shown in Fig. 2 (b). Task 3 explores lesion tracking, where we register the moving image with the lesion to the fixed image without the lesion. This task is designed to evaluate whether the registration model attempts to erase or suppress the lesion in the moving image in order to match the target. For Task 4, we register PET-CT to MRI, which is more challenging due to both the cross-modality nature and the significant breast deformation caused by different patient postures (lying on the back versus facing down).
@@
Evaluation metrics: Tasks 1 to Task 3 are evaluated using the Dice Similarity Coefficient ($DSC$) of three targets: breast, FGT, and organ. For Task 3, we further assess lesion preservation by calculating $V_{\varepsilon}=abs(\sum M_{les-warped}-\sum M_{les-mov})/\sum M_{les-mov}$, which reflects the change in lesion size during registration. For lesion tracking purposes, the ideal goal is to align the breast and dense tissue while preserving the lesion’s physical characteristics, rather than altering its shape or size during the registration process. For Task 5, we qualitatively evaluate the registration performance by visualizing the warped CT and overlaying the PET onto the fixed MRI.
@@ 4 Results and Discussion
Figure 2: (a). Visualization of example registration results for different foundation model-based algorithms. Dense tissue from the fixed image slice is shown in green, and the warped dense tissue is shown in orange. The overlapping regions appear in yellow; (b). subgroup analysis of the registration performance versus the sequence of the moving image; (c)-(d). Visualization results of selected cases for task 3 and task 4.
Table 1 shows that foundation models, especially SAM, achieve superior performance in large-structure registration like breast contour. When registering volumes from the same MRI sequence from longitudinal exams, the performance of foundation-based models and optimization-based methods is comparable. However, for cross-sequence registration, SAM noticeably outperforms the best optimization-based method, achieving a Dice score (DSC) of $80.84\%$ versus $76.82\%$. This suggests that features extracted by foundation models are domain-invariant, making them more robust under shifts in image appearance. In contrast, optimization-based methods rely more heavily on image-level similarity.
However, we also observe that foundation model-based algorithms show worse performance in FGT registration for both Task 1 and Task 3, and similar or even worse performance in Task 2. To better understand this, we visualize the feature maps extracted by each model in Fig. 3. The results show that FGT details are barely visible in these feature maps, especially in sparse FGT cases, such as the example in row 3. This limitation is also reflected in the qualitative results in Fig. 2 (a), where, although most models (except MedCLIP-SAM) can register the overall breast contour well, FGT misalignment remains substantial.
For organ registration, although we initially hypothesized that models pre-trained on medical-domain data would perform better due to exposure to similar anatomical structures, all foundation models performed worse than optimization-based methods. Regarding Task 3 (lesion tracking), DINO-v2 achieves the best performance in preserving lesion size, while MedSAM performs poorly.
@@
For Task 4 (PET-CT to MRI registration), the benefits of foundation models become more evident. As shown in Fig. 2 (d), NiftyReg completely fails to align organs between CT and MRI, SAM successfully registers CT to MRI despite significant shape differences, confirming the advantage of foundation models under large domain gaps (e.g., CT to MRI) [26]. This result also suggests potential applications in overlaying PET onto MRI to visualize tumors.
When comparing models trained on natural images versus those trained on medical datasets (such as MedSAM) or breast MRIs (SSLSAM), we observe mixed results. For SSLSAM, there is no clear improvement in Task 1 and a performance drop in Task 2. We suspect this may be due to SSLSAM’s pre-training on only pre-contrast MRIs, leading to domain overfitting. For MedSAM, we observe substantial performance drops across all tasks and targets, likely due to pre-training on medical datasets that excluded breast MRIs. MedCLIP-SAM fails across all tasks, possibly due to the low resolution of feature maps (Fig. 3).
@@
Conclusion. In this study, we explored whether foundation models can be used for zero-shot registration tasks in the context of breast MRI. We find that models pre-trained on natural images, such as SAM and DINO-v2, can achieve strong performance for large-structure alignment, particularly for breast contour registration. However, these models struggle to preserve fine anatomical details, such as FGT, limiting their application in tasks that require accurate tracking of internal structures. We believe this highlights an important direction for future research, which could develop strategies to better preserve fine-grained information within the feature representations of foundation models.
Appendix 0.A Implementation details for each foundation model
For SAM-related models, including SAM, MedSAM, and SSL-SAM, the input slices are resized to $1024\times 1024$. For DINO-v2, the input size is flexible, so we upsample the slices to the largest size that can fit into a single GPU A6000, which is $1792\times 1792$, ensuring the size is divisible by 14 to match the patch size. For MedCLIP-SAM, the input slices are resized to $224\times 224$. All inputs are normalized according to the pretraining settings of each model, using the normalization methods provided in their respective codebases. For the dimension reduction step, the size of the extracted feature map for each slice ($F^{z}$) varies by model, as described in Fig 3. Based on this, we set the number of retained principal components to $d=3$ for SAM-based models, $d=8$ for DINO-v2, and $d=12$ for MedCLIP-SAM. The choice of $d$ considers two factors: (1) models with higher original feature dimension $D$ require more components to preserve information, and (2) hyperparameters are optimized on one selected reference image pair. For the registration step, we follow the hyperparameters selected by DINO-reg [26].
Figure 3: (a) Left: characteristics of each foundation model’s encoder, including their training strategy and dataset domain; (b) Right: examples of the first three channels of the dimension-reduced feature map by each model, visualized as RGB.
Appendix 0.B Baseline Methods to Compare
In this study, we provide four optimization-based registration algorithms as baseline methods to evaluate the performance of foundation-model-based approaches. Specifically, we include (1) affine registration implemented using NiftyReg [20] to account for global linear transformations, (2) the Symmetric Normalization (SyN) [2] algorithm from ANTs [28], a widely adopted deformable registration method known for its accuracy in capturing complex non-linear deformations, and (3) NiftyReg-based deformable registration [21], which leverages free-form deformation using cubic B-splines to model local anatomical variations. (4) ConvexAdam MIND [25], compared to other optimization-based registration algorithms, replaces commonly used similarity metrics such as normalized cross correlation (NCC) and mutual information (MI) with the Modality Independent Neighborhood Descriptor (MIND), which captures local structural information and is robust to intensity variations across modalities. The implementation details for these four methods are presented below, respectively.
Affine: For Affine registration, we utilize the reg_aladin tool from the NiftyReg [20] library to align the moving image to the fixed image using a global linear transformation. The resulting affine matrix is applied to breast, dense tissue, and organ masks with nearest-neighbor interpolation.
SyN: SyN-based [2] deformable registration baseline is performed utilizing the antsRegistration tool from the ANTs [28] library. The gradient step size is 0.1, with a total field variance regularization of 3, and mutual information (MI) is used as the similarity metric. A multi-resolution optimization strategy is adopted with four levels of resolution, defined by shrink factors 8 $\times$ 4 $\times$ 2 $\times$ 1 and smoothing sigmas 3 $\times$ 2 $\times$ 1 $\times$ 0. The convergence criteria are set as [100x70x50x20, 1e-6, 10], indicating the maximum number of iterations per level, the convergence threshold, and the number of consecutive steps below the threshold required for convergence.
NiftyReg: reg_f3d tool from the NiftyReg library is utilized to perform the NiftyReg-based deformable registration [21]. We utilize the default hyperparameters in the reg_f3d implementation.
ConvexAdam MIND: The implementation of ConvexAdam follows the official implementation provided by Siebert et al. [25]. For feature extraction, we follow MIND with the Self-Similarity Context (SSC) extension. The neibohood size used for calculating similarity is set as $radius=3$ and the dialation size is 2.
Appendix 0.C Exam Collection
For the MRI-to-MRI registration tasks, the data is collected internally from our institution, consisting of exams performed under the bilateral breast MRI protocol, ranging from the year 2014 to the year 2021. We only include patients who have multiple time-stamped exams and at least one pre-contrast axial view volume available. For the PET-CT to MRI registration task, the data is collected externally from BREAST-DIAGNOSIS collection [4], where we select patients who have both PET-CT and MRI scans acquired within a six-month interval. All exams are first resampled into $1mm\times 1mm\times 1mm$ isotropic voxel spacing and then padded into $512\times 512\times 512$ as volume size. For each task, we randomly selected fifty pairs from the pairs that meet the task criteria, with the task details described in a later paragraph Tasks. From these pairs, volumes from patients who had undergone mastectomy were excluded.
Appendix 0.D Mask Collection
We include three targets for our registration evaluation: breast, breast fibroglandular tissue (FGT), and organ. The first two, breast and FGT, are important targets to register for breast-specific tasks, as accurate alignment of both overall breast shape and internal dense tissue is critical for clinical and research use. For the organ label, we suspect that pre-trained networks may perform relatively well, as this type of structure is commonly seen in abdominal MRI or CT datasets that many of these models were originally exposed to. For breast and FGT mask, we employ the breast segmentation model given by the previous study utilizing 3D V-Net architecture [7, 16] followed by the manual correction. The organ and lesion masks are generated using a publicly available breast MRI segmentation model based on nnU-Net, followed by manual correction. The generated masks are visualized in the right-hand side of Fig. 3, in the Mask column.

### v2

@@ Abstract
Vision foundation models (VFM), pre-trained on large image datasets and capable of capturing rich feature representations, have recently shown potential for zero-shot image registration. However, their performance has mostly been tested in the context of less complex structures, such as the brain or abdominal organs, and it remains unclear whether these models can handle more challenging, deformable anatomy. Breast MRI registration is particularly challenging due to anatomical variation between patients, deformation from positioning, and the thin, complex structure of fibroglandular tissue (FGT), where accurate alignment is crucial. In this study, we provide a comprehensive evaluation of VFM-based registration algorithms. We assess five pre-trained encoders, including DINO-v2, SAM, MedSAM, SSLSAM, and MedCLIP, across four key breast registration tasks that capture variations in different years and dates, sequences, modalities, and patient disease status (lesion versus no lesion). Our results show that among the five pre-trained encoders, SAM demonstrates a clear advantage. VFM outperforms traditional optimization-based methods, especially in cross-modality tasks, with its advantage more pronounced for large structures (e.g., organs and breast) than for finer anatomical details like FGT. Further work is needed to understand how domain-specific training influences registration and to explore targeted strategies that improve fine structure accuracy. We also publicly release our code at Github.
@@
Foundation model Breast registration Zero-shot.
@@ 1 Introduction
Figure 1: Pipeline for VFM-based Breast Image Registration. The Found Encoder can be replaced with any pretrained vision foundation encoder.
In contrast, vision foundation models (VFMs) have gained attention for their strong generalization ability and zero-shot performance on downstream tasks, such as segmentation and classification [22, 18, 9]. One recent effort, DINO-Reg [24], shows that DINO-v2’s rich semantic embeddings can support effective zero-shot registration across medical imaging domains. Yet, it remains unclear how different pre-training strategies, such as masked autoencoders (MAE), vision-language models (VLMs), or medical domain-specific models like MedSAM, affect registration performance. Moreover, most studies focus on less deformable regions like abdominal organs with larger, more stable structures, while breast registration poses unique challenges with significant shape variation and thin, highly deformable tissue. How VFM-based algorithms handle such complex anatomy remains poorly understood.
@@
    We conduct a comprehensive comparison of registration strategies according to the pretraining strategies that were applied to develop the foundation model, including contrastive learning, masked autoencoder, and vision-language approaches, across both natural and medical imaging domains.
@@
    We design four registration tasks that reflect real-world challenges, covering longitudinal, cross-sequence, lesion-tracking, and PET-to-MRI alignment.
@@
    We systematically evaluate the strengths and limitations of VFM-based algorithms compared to other zero-shot approaches.
@@ 2 Method
2.1 VFM-based registration
Our training-free, VFM-based registration pipeline, inspired by the framework of  [24], includes three main stages: (1) feature extraction, (2) dimension reduction, and (3) feature-based registration, as shown in Fig. 1.
@@ 2.1 VFM-based registration
To prepare for volumetric registration, we reshape $\tilde{F}_{\text{ref}}^{z}$, $\tilde{F}_{\text{mov}}^{z}$ back to a 3D spatial format. Specifically, each slice-level feature map $\tilde{F}^{z}\in\mathbb{R}^{N\times d}$ is first reshaped into a 2D grid of size $w\times h\times d$, where $w$ and $h$ represent the number of patch tokens along the width and height, respectively. Stacking the reshaped slices along the $z$-axis results in a 4D volume. This intermediate volume is then upsampled to match the original spatial resolution, resulting in a final feature volume of size $W\times H\times Z\times d$, which aligns with the input dimensions and serves as input to the registration step.
@@
[…]
@@ 2.2 Foundation models
SAM [13]: Trained on large-scale natural image segmentation datasets using a promptable segmentation objective. The model adopts a ViT-based encoder and was first pre-trained with MAE, followed by additional prompt-based segmentation training. SAM has also been widely used for general vision tasks. Although SAM is designed for prompt-based segmentation, we use only its image encoder to automatically generate latent embeddings without human input.
MedSAM [17]: A domain-specific version of SAM, fine-tuned on various medical image datasets including MRIs under a prompt-based setting to improve performance on medical image segmentation tasks. Notably, as shown in the original MedSAM Supplementary Tables 1–4, the training datasets did not include any breast MRI or breast CT images.
SSLSAM [10]: Fine-tuned from SAM’s initial weights on various MRI datasets without labels under the MAE setting. Noticeably, their pre-trained dataset contains one publicly available breast dataset for dynamic pre-contrast MRIs, but not other breast MRI sequences, e.g., post-contrast [16].
@@
Figure 2: (a) Encoder characteristics for each VFM (training strategy and dataset domain). (b) Ablation of dimensionality reduction (PCA dimension $d$) versus registration performance for each encoder. (c) First three channels of the dimension-reduced feature map, visualized as RGB. (d) Top feature channel from raw feature map showing the highest FGT contrast, aiming to visualize whether FGT-related information is captured in the latent features.
@@ 3.1 Datasets preparation
Exam collection. For the MRI-to-MRI registration tasks, the data is collected internally from our institution, consisting of exams performed under the bilateral breast MRI protocol, ranging from the year 2014 to the year 2021. We only include patients who have multiple time-stamped exams and at least one pre-contrast axial view volume available. For the PET-CT to MRI registration task, the data is collected externally from BREAST-DIAGNOSIS collection [4], where we select patients who have both PET-CT and MRI scans acquired within a six-month interval. All exams are first resampled into $1mm\times 1mm\times 1mm$ isotropic voxel spacing and then padded into $\times 512\times 512$ 512 as volume size. For each task, we randomly selected fifty pairs from the pairs that meet the task criteria, with the task details described in a later paragraph Tasks. From these pairs, volumes from patients who had undergone mastectomy were excluded.
Mask collection. We include three targets for our registration evaluation: breast, FGT, and organ. The first two, breast and FGT, are important targets to register for breast-specific tasks, as accurate alignment of both overall breast shape and internal dense tissue is critical for clinical and research use. For the organ label, we suspect that pre-trained networks may perform relatively well, as this type of structure is commonly seen in abdominal MRI or CT datasets that many of these models were originally exposed to. For breast and FGT mask, we employ the breast segmentation model given by the previous study utilizing 3D V-Net architecture [7, 16] followed by the manual correction. The organ and lesion masks are generated using a publicly available breast MRI segmentation model based on nnU-Net, followed by manual correction. The generated masks are visualized in the right-hand side of Fig. 2, in the Mask column.
@@
Tasks design. In this study, we design four tasks and one ablation study to comprehensively evaluate the registration performance of multiple models. Task 1 involves registering breast MRI scans from different dates and years (longitudinal exams) with the same image sequence (all images are dynamic pre-contrast). Task 2 focuses on registering longitudinal breast MRI exams with different image sequences, where the fixed images are set to the dynamic pre-contrast sequence. The category of moving image sequences is shown in Fig. 3 (b). Task 3 explores lesion tracking, where we register the moving image with the lesion to the fixed image without the lesion. This task is designed to evaluate whether the registration model attempts to erase or suppress the lesion in the moving image in order to match the target. For Task 4, we register PET-CT to MRI, which is more challenging due to both the cross-modality nature and the significant breast deformation caused by different patient postures (lying on the back versus facing down). We also include the ablation study to disentangle the effects of feature representation and optimization strategy, we replaced ConvexAdam with standard registration methods, SyN and NiftyReg FFD, to register SAM-derived feature volumes in three randomly selected volume pairs from Task 2. The result for all tasks are shown in Table 1.
@@
Evaluation metrics. Tasks 1 to Task 3 are evaluated using the Dice Similarity Coefficient ($DSC$) of three targets: breast, FGT, and organ. For Task 3, we further assess lesion preservation by calculating $V_{\varepsilon}=abs(\sum M_{les-warped}-\sum M_{les-mov})/\sum M_{les-mov}$, which reflects the change in lesion size during registration. For lesion tracking purposes, the ideal goal is to align the breast and dense tissue while preserving the lesion’s physical characteristics, rather than altering its shape or size during the registration process. For Task 4, we qualitatively evaluate the registration performance by visualizing the warped CT and overlaying the PET onto the fixed MRI.
3.2 Implementation details of VFM-based registration
For SAM-related models, including SAM, MedSAM, and SSL-SAM, the input slices are resized to $\times 1024$ 1024. For DINO-v2, the input size is flexible, so we upsample the slices to the largest size that can fit into a single GPU A6000, which is $\times 1792$ 1792. For MedCLIP-SAM, the input slices are resized to $\times 224$ 224. All inputs are normalized according to the pretraining settings of each model, using the normalization methods provided in their respective codebases. For the dimension reduction step, the size of the extracted feature map for each slice ($F^{z}$) varies by model. The PCA dimension $d$ for each encoder is optimized on 9 sampled pairs from Task 3. We select the $d$ at which DSC plateaus, balancing performance and computational cost, as higher $d$ increases memory usage and slows registration (Fig. 2). Based on this, we set $d=16$ for SAM, $d=10$ for MedSAM, $d=16$ for SSLSAM, $d=10$ for DINO-v2, and $d=16$ for MedCLIP-SAM. The choice of $d$ is based on an ablation from $d=1$ to 161616 (capped at $d=12$ for DINO-v2 due to GPU limits). For the registration step, we follow the hyperparameters selected by DINO-reg [24].
@@ 4 Results and Discussion
Ablation study: contribution of features vs. optimizer
Raw + SyN 71.1 $\pm$ 30.4 40.9 $\pm$ 18.5 65.7 $\pm$ 30.0 -
Raw + NiftyReg 70.5 $\pm$ 20.4 50.5 $\pm$ 19.0 70.7 $\pm$ 6.5 -
Raw + ConvexAdam 69.4 $\pm$ 25.4 46.2 $\pm$ 22.7 63.7 $\pm$ 11.3 -
SAM + SyN 58.9 $\pm$ 32.7 24.7 $\pm$ 11.1 53.7 $\pm$ 24.8 -
SAM + NiftyReg 74.9 $\pm$ 26.5 35.7 $\pm$ 20.7 69.7 $\pm$ 10.7 -
SAM + ConvexAdam 85.0 $\pm$ 18.8 47.0 $\pm$ 27.9 73.4 $\pm$ 9.2 -
[…]
@@
Regarding Task 3 (lesion tracking), DINO-v2 achieves the best performance in preserving lesion size, while MedCLIP-SAM performs poorly. For Task 4 (PET-CT to MRI registration), the benefits of foundation models become more evident. As shown in Fig. 3 (d), NiftyReg completely fails to align organs between CT and MRI, SAM successfully registers CT to MRI despite significant shape differences, confirming the advantage of foundation models under large domain gaps (e.g., CT to MRI) [24]. This result also suggests potential applications in overlaying PET onto MRI to visualize tumors.
Discussion for pre-train strategies. Although both SAM and DINO-v2 are pre-trained on natural images, we observe a performance gap here. We attribute this to differences in their training strategies: SAM uses a two-stage pipeline, where the second stage is explicitly trained for pixel-wise segmentation, enabling it to learn fine-grained, spatially localized, and structure-aware representations. DINO-v2, on the other hand, is trained via contrastive learning. While it includes local and global tokens, the patch sizes are relatively large, limiting its ability to capture detailed spatial information. Compared to MedCLIP-SAM, the feature tokens are even coarser, and the model is originally designed for image-text alignment rather than local-information preserving. To further understand these differences, we visualize the feature maps from each model in Fig.2. For row 3, only SAM and SSLSAM capture some FGT structures, while others fail. Even for SAM/SSLSAM, FGT details is barely visible in the feature map, this limitation is also evident in the qualitative results (Fig. 3 (a)), where models generally register the breast contour well, but substantial FGT misalignment remains. The zero-shot performance of the best VFM-based method still lags significantly behind other training-based approaches, such as GuidedMorph [6].
Discussion for pre-train datasets. When comparing models trained on natural images versus those trained on medical datasets (e.g., MedSAM) or pre-contrast breast MRIs (SSLSAM), we observe mixed results. SSLSAM shows improvement in FGT and organ registration in Task 1, but performance drops in Task 2. We suspect this is due to SSLSAM’s pretraining solely on pre-contrast breast MRIs [10], leading to domain overfitting. As a result, while SSLSAM performs reasonably well for pre-contrast MRI alignment (Task 1), it performs worse other unseen sequences in Task 2. This is also evident in Fig. 3(b), where SSLSAM performs better on sequences similar to its pretraining data but shows noticeably worse performance on others, such as SUB. For MedSAM, we observe substantial performance drops across all tasks for breast and FGT. This likely results from its pretraining on a limited set of medical imaging datasets (approximately 1 million images, compared to 11 million natural images used for SAM), none of which included breast MRIs or CT, as indicated in the original MedSAM paper. However, it contains substantial data for abdominal MRIs, which makes it capable of capturing organ features effectively. As shown in Fig. 2, its feature maps primarily emphasize organ structures. MedCLIP-SAM fails across all tasks despite being trained on MRIs. Due to the low resolution of its feature maps (Fig. 2), it is far from suitable for registration.
@@
Discussion about ablation study. We conducted two groups of ablation studies. The first explores the effect of PCA dimension ($d$) on registration performance, as shown in Fig. 2. We observed that when $d$ is too low, it may discard important anatomical details necessary for accurate registration, as illustrated by the first three dimensions in Fig. 2 (c). However, as $d$ increases, the performance plateaus and reaches a relatively stable level. Since larger $d$ values significantly increase memory usage and computation time, we find that selecting $d$ in the range of 10–16 offers a good trade-off for each encoder. The second study evaluates the contribution of feature maps and optimizers to registration performance. As shown in Table 1, SAM’s extracted features outperform raw images, and the combination of SAM features with the ConvexAdam optimizer achieves the best performance among all VFM-based registration methods.
Figure 3: (a). Visualization of example results for different VFM-based algorithms. FGT from the fixed image slice is shown in blue, and the warped dense tissue is shown in purple. The overlapping regions appear in green; (b). subgroup analysis of the registration performance versus the sequence of the moving image; (c)-(d). Visualization results of selected cases for task 3 and task 4.
@@
Conclusion. In this study, we explored whether foundation models can be used for zero-shot registration tasks in the context of breast MRI. We find that models pre-trained on natural images, such as SAM, can achieve strong performance for large-structure alignment, particularly for breast contour registration. However, these models still struggle to preserve fine anatomical details, such as FGT, limiting their application in tasks that require accurate tracking of internal structures. We believe this highlights an important direction for future research, which could develop strategies to better preserve fine-grained information within the feature representations of foundation models.
Acknowledgments The research reported in this publication was supported by the National Institute of Biomedical Imaging and Bioengineering of the National Institutes of Health under the award number R01EB031575. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.